OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2

OpenStudy (anonymous):

A.1 B.0 C.2 D.infinitely many

OpenStudy (amoodarya):

multiply 2nd by 3 then sum with the first equation simplify then you have value of x

OpenStudy (anonymous):

2nd by 3?

OpenStudy (amoodarya):

yes multiply 2nd by 3

OpenStudy (anonymous):

the second equation by 3?

OpenStudy (amoodarya):

3( 2x-y=2) 6x-3y=6

OpenStudy (amoodarya):

and now 1st +2nd = ?

OpenStudy (anonymous):

so the 2nd equation is now 6x-y=6?

OpenStudy (anonymous):

so I would get 12x-4y=12

OpenStudy (amoodarya):

5x + 3y = 7 6x -3y =6 sum of them is 11x=7+6

OpenStudy (anonymous):

oh ok

OpenStudy (amoodarya):

so ?

OpenStudy (anonymous):

so x=1?

OpenStudy (anonymous):

wall its more x=1.6

OpenStudy (anonymous):

*well

OpenStudy (amoodarya):

x=13/11 put x to one of equation so you got y now you have one solution for the system of equation

OpenStudy (amoodarya):

do find out ?

OpenStudy (amoodarya):

sorry i meant do you find out ?

OpenStudy (anonymous):

yes y=1?

OpenStudy (anonymous):

y=-5/3x+7/3

OpenStudy (amoodarya):

no y=4/11

OpenStudy (amoodarya):

and x=13/11

OpenStudy (anonymous):

so the answer is D?

OpenStudy (anonymous):

infinitely many

OpenStudy (amoodarya):

no the system has 1 solution it is (x=13/11 ,y=4/11)

OpenStudy (anonymous):

thanks